# Comparative Analysis of $k$-essence and Quintessence Scalar Field Models: A Data Analysis Approach

### v1

@@ Abstract
We perform a comparative analysis of quintessence and $k$-essence scalar field models in the data analysis perspective. We study the quintessence field with an exponential potential and the $k$-essence field with an inverse square potential in the present work. Before delving into data analysis, we provide a brief perspective on dynamical evolution on both of the models and obtain the stability constraints on the model parameters. We adopt Bayesian inference procedure to estimate the model parameters that best-fit the data. A comprehensive analysis utilizing Observational Hubble data (OHD) and Pantheon+ compilation of Type Ia supernovae (SNIa) shows that $k$-essence model fits the data slightly better than the quintessence model while the evidence of these models in comparison with the $\Lambda$CDM model is weak. The value of the Hubble constant predicted by both the models is in close agreement with the value obtained by the Planck2018 collaboration assuming the $\Lambda$CDM model.
@@ I Introduction
Over the past two decades, researchers have explored various alternative scenarios that can mimic the dynamics of $\Lambda$CDM. One of the most popular and heavily investigated mechanisms is the introduction of a scalar field. Inspired by particle physics, a quintessence-type field with a potential term can generate sufficient acceleration in the presence of dark matter [19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. Although the quintessence field minimally coupled with dark matter fluid can mimic the dynamical aspect of $\Lambda$CDM, it also suffers from fine-tuning of the potential parameter. Later, some string-inspired models, namely $k$-essence fields with non-linear kinetic terms, have been widely studied, as they were claimed to address some of the shortcomings of the above models. Due to the associated non-linear kinetic terms, these fields have been termed non-canonical fields [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. Despite its success in explaining late-time cosmic acceleration, few attempts have been made to test these models against observational data due to the complexity of solving the Hubble equation to obtain an analytical expression for the Hubble parameter.
It is essential to test the $k$-essence model against observational data to determine its compatibility. To conduct this analysis, we solve the field and Friedmann equations numerically and obtain the Hubble parameter as a function of redshift. To set initial conditions and priors for the model parameters, we perform a dynamical system analysis. Stability techniques in dynamical systems are powerful tools for studying complex models, as they help identify physically viable solutions by examining the existence and nature of critical points, thereby constraining the model parameters [40, 41, 42, 43, 44, 45, 46, 47]. To analyze the stability of the critical points, we set up a system of first-order autonomous differential equations. By linearizing these autonomous equations, we construct a Jacobian matrix and obtain the corresponding eigenvalues. If any real part of the eigenvalues vanish, the linearization method is no longer applicable for determining stability in those cases. In such situations, more rigorous mathematical techniques, such as the center manifold theorem, must be utilized.
In this study, we compare the quintessence and $k$-essence fields, using exponential and inverse power-law type potentials, respectively, against observational Hubble data (OHD) and Pantheon+ datasets. The dynamics of the quintessence field are potential-dependent, and this field has been extensively studied with a wide range of potentials. [48, 49, 50]. However, very few studies have been conducted on purely kinetic $k$-essence, and $k$-essence with a Lagrangian which is a general function of $X$ and the $\phi$ where $\phi$ is the $k$-essence field and $X=-1/2g^{\mu\nu}\nabla_{\mu}\phi\nabla_{\nu}\phi$ [51, 52]. For the first time, we constrain the parameter space of the $k$-essence field, where the Lagrangian takes the form $\mathcal{L}_{\phi}=V(\phi)F(X)$ and compare it with the $\Lambda$CDM model in the data analysis perspective.
The paper is systematized as follows. In section II, we present quintessence model and its dynamical system analysis. In section III, we analyse the $k$-essence model in the dynamical system perspective. Data analysis of the quintessence and $k$-essence models are presented in section IV. Finally, we conclude in section V. The model comparison performed based on the AIC and BIC criteria shows that the $k$-essence model is slightly better than the quintessence model.
II The quintessence field
The action describing the quintessence field minimally coupled with the dark matter fluid is given by:
$S=\int d^{4}x\sqrt{-g}\bigg{(}\frac{R}{2\kappa^{2}}+\mathcal{L}_{\phi}(\partial{}_{\mu}\phi,\phi)+\mathcal{L}_{m}\bigg{)},$ (1)
where,
$\mathcal{L}_{\phi}=\left[\epsilon\dfrac{1}{2}\partial{}_{\mu}\phi\partial{}^{\mu}\phi-V(\phi)\right]\,.$ (2)
Here, $\epsilon$ serves as a switching parameter with values of ±$\pm 1$. $\epsilon=+1$ represents the phantom field, while $\epsilon=-1$ represents the normal quintessence field. In our system of units $\kappa^{2}=8\pi G$ where $G$ is Newton’s universal gravitational constant.
Variation of the action with respect to $\phi$ and the metric $g^{\mu\nu}$ yields the field equations and the energy-momentum tensor of the field:
$\begin{split}\epsilon\nabla_{\mu}(\nabla^{\mu}\phi)+\frac{dV}{d\phi}&=0,\\ T_{\mu\nu}^{(\phi)}&=-\epsilon\partial{}_{\mu}\phi\partial{}_{\nu}\phi+g_{\mu\nu}\left[\frac{\epsilon}{2}\partial{}_{\alpha}\phi\partial{}^{\alpha}\phi-V(\phi)\right]\,.\end{split}$ (3)
The corresponding energy density and pressure are given by:
$\rho_{\phi}=-\frac{\epsilon\dot{\phi}^{2}}{2}+V(\phi),\quad P_{\phi}=-\frac{\epsilon\dot{\phi}^{2}}{2}-V(\phi)\,.$ (4)
The equation of state of the field is expressed as:
$\omega_{\phi}=\frac{-\frac{\epsilon\dot{\phi}^{2}}{2}-V(\phi)}{-\frac{\epsilon\dot{\phi}^{2}}{2}+V(\phi)}.$ (5)
For the normal field, to achieve cosmic acceleration, it is required that $V(\phi)\gg\dot{\phi}^{2}/2$. The Friedmann equation for a spatially flat Friedmann-Lemaitre-Robertson-Walker (FLRW) metric, $ds^{2}=-dt^{2}+a(t)^{2}d\vec{x}^{2}$, is given by
$\begin{split}3H^{2}&=\kappa^{2}\left(\rho_{m}+\rho_{\phi}\right),\\ 2\dot{H}+3H^{2}&=-\kappa^{2}\left(P_{m}+P_{\phi}\right),\end{split}$ (6)
where $H\equiv\dot{a}/a$ is the Hubble parameter and $a(t)$ is the scale-factor of the expanding universe. Here, $\rho_{m}$ and $\rho_{\phi}$ represent the energy densities of matter and field components, respectively, $P_{m\,,\,\phi}$ denotes the pressure of the respective components. The energy-momentum tensor of both the field and fluid are individually conserved, resulting in the following conservation equations:
Note that during the late-time epoch, the universe is primarily dominated by pressureless dust and dark energy. Consequently, one can safely neglect the contribution of radiation from the system. Quintessence has been extensively studied with a wide variety of potentials, ranging from power-law types to exponential potentials. In this paper, we will specifically focus on demonstrating the dynamics for an exponential-type potential [23],
$V(\phi)=V_{0}e^{-\lambda\kappa\phi}\,,$ (8)
where $V_{0}$ and $\lambda$ are constants. In the subsequent section, we will set up the dynamical equations within the framework of dynamical system analysis to understand the stability and nature of solutions.
II.1 Dynamical system for quintessence fields
To understand the dynamics, we begin by constructing the autonomous system of equations through the definition of dimensionless variables as follows:
$x$ $=\frac{\kappa\dot{\phi}}{\sqrt{6}H},\quad y=\frac{\kappa\sqrt{V}}{\sqrt{3}H},\quad\lambda=-\frac{V_{,\phi}}{\kappa V},\quad\Omega_{m}={\dfrac{\kappa^{2}\rho_{m}}{3H^{2}}},\quad\Omega_{\phi}=\frac{\kappa^{2}\rho_{\phi}}{3H^{2}}.$ (9)
These are the standard definitions of the variables commonly adopted in the literature to describe the dynamics of the system. The parameter $x$ is related to the dynamic velocity of the field, while $y$ describes the evolution of the potential. The fractional densities of the fluid and field are represented by $(\Omega_{m},\Omega_{\phi})$ respectively, and $\lambda$ is related to the slope of the potential. For the exponential potential, $\lambda$ becomes a constant. A crucial constraint on the system is imposed by the first Friedmann equation Eq. (6),
$1=\Omega_{m}+\Omega_{\phi},$ (10)
where the physically viable solution must adhere to these limits,
$0\leq\Omega_{m}\leq 1,\quad 0\leq\Omega_{\phi}\leq 1$ (11)
Here, the field fractional density and effective equation of state (EoS) for the quintessence field $(\epsilon=-1)$ is given by:
$\Omega_{\phi}=x^{2}+y^{2},\quad\omega_{\rm eff}=\frac{P_{m}+P_{\phi}}{\rho_{m}+\rho_{\phi}}=x^{2}-y^{2}$ (12)
and the second Friedmann equation, Eq. (6), can be expressed in terms of dynamical variables as:
$\frac{2\dot{H}}{3H^{2}}=-\left(1+\omega_{m}\Omega_{m}+x^{2}-y^{2}\right)\,,$ (13)
where $\omega_{m}$ is the EoS in the matter sector. The autonomous system of equations are as follows:
$\begin{split}x^{\prime}&=\left(\dfrac{3y^{2}\lambda}{\sqrt{6}}-3x\right)+\dfrac{3}{2}x\left(1+\omega_{m}\Omega_{m}+x^{2}-y^{2}\right),\\ y^{\prime}&=\dfrac{-xy\lambda\sqrt{6}}{2}+\dfrac{3y}{2}(1+\omega_{m}\Omega_{m}+x^{2}-y^{2}).\end{split}$ (14)
Here, prime denote derivative with respect to N, where $dN=Hdt$. Since the autonomous equations of the system depend only on the two variables of the system, the phase space becomes 2D. The dynamics of the matter fluid can be understood using the Friedmann constraint. The autonomous equation $(y^{\prime})$ is invariant under $(y\mapsto-y)$, and has an invariant submanifold at $y=0$. Consequently, trajectories originating on this plane remain confined to it. The critical points and their eigenvalues are summarized in Tab. [1] by choosing $\omega_{m}=0$. As the matter sector is assumed to be made up of dark matter, our assumption implies that the fluid representing dark matter is pressure-less. We use linearization techniques to determine the stability of the critical points by evaluating the Jacobian matrix defined as follows:
$J_{ij}=\frac{\partial x_{i}^{\prime}}{\partial x_{j}}\bigg{|}_{\bm{x}_{*}},$ (15)
where $x_{i}=\{x,y\}$ for $i=1,2$. Corresponding to this autonomous system, the Jacobian matrix becomes $2\times 2$ dimensional. To assess stability, we determine the eigenvalues and based on the signature of the real parts of eigenvalues, the point can be categorized as stable, unstable, or saddle. If all the real parts of the eigenvalues are negative (positive), the corresponding point becomes asymptotically stable (unstable). However, for alternating signs, the point is labeled as a saddle.
Points $(x_{*},y_{*})$ $\Omega_{\phi}$ $\Omega_{m}$ $\omega_{\rm eff}$ Eigenvalues ($E_{1},E_{2}$)
$P_{1,2}$ $(\mp 1,0)$ 1 0 1 $\left(3,3-\sqrt{\frac{3}{2}}\lambda\right)$
$P_{3}$ $\left(\frac{\sqrt{\frac{3}{2}}}{\lambda},\frac{\sqrt{\frac{3}{2}}}{\lambda}\right)$ $\frac{3}{\lambda^{2}}$ $1-\frac{3}{\lambda^{2}}$ 0 $\left(-\frac{3\left(\lambda^{2}+\sqrt{24\lambda^{2}-7\lambda^{4}}\right)}{4\lambda^{2}},\ \frac{3\sqrt{24\lambda^{2}-7\lambda^{4}}}{4\lambda^{2}}-\frac{3}{4}\right)$
$P_{4}$ $\left(\frac{\lambda}{\sqrt{6}},\sqrt{1-\frac{\lambda^{2}}{6}}\right)$ 1 0 $\frac{1}{3}\left(\lambda^{2}-3\right)$ $\left(\frac{1}{2}\left(\lambda^{2}-6\right),\lambda^{2}-3\right)$
$P_{5}$ $(0,0)$ 00 1 0 $\left(-\frac{3}{2},\frac{3}{2}\right)$
Table 1: The critical points of minimally coupled quintessence field.
The points $P_{1,2}$ specify points where the system is kinetically drive. At this points the potential term vanishes and consequently $y=0$. Near this points the effective EoS is 1. The other point $P_{5}$, where both field parameters vanish, results in fluid density domination over the field. As a consequence, the effective $\omega_{\rm eff}$ becomes zero, signifying matter-dominated characteristics. This point always exhibits saddle-type behavior.
The point $P_{3}$ is dependent on the model parameter, and the corresponding equation of state is zero, which may signify the matter regime of the universe. However, from the Friedmann constraint, Eq. (10), the point becomes nonphysical for $\lambda^{2}\leq 3$. The point can only stabilize for $3\leq\lambda^{2}\leq\frac{24}{7}$. Beyond this range, it becomes unstable.
The point $P_{4}$ is also model parameter dependent. At this point, the field density dominates over the fluid density and the equation of state becomes $\lambda$ dependent. For $\lambda^{2}=3$, the point can generate non-relativistic matter type behavior as $\omega_{\rm eff}$ vanishes. However, this point can also produce an accelerating solution for $20<\lambda^{2}<$ 2. It is clear that if we select the model parameter in this range, the point $P_{3}$ becomes non-physical. The point $P_{4}$ can stabilize for $0\leq\lambda^{2}\leq 3$. This shows that at $\lambda=\sqrt{3}$, the quintessence field behaves as a non-accelerating attractor type field.
We assess the dynamics of the entire system in Fig.[1] for the range of $0.001\leq x_{0}\leq 0.2$, $0.7\leq y_{0}\leq 0.95$, $0.0001\leq\lambda\leq 1.3$ by generating random numbers within this range. Here, the subscript zero denotes the initial conditions at $N=z=0$. We have also plotted the evolution of the Hubble parameter and distance modulus, $\mu$, against redshift, $z$, where $\mu$ is defined as
$\mu=5\log_{10}(d_{L})+25,$ (16)
where the luminosity distance is given by:
[…]
for the random numbers generated within the parameter ranges specified above. Here, $c$ represents the speed of light expressed in km/s. We compare it with the corresponding observational data including 43 OHD and 1701 Pantheon+ data sets [53, 54].
From this evolution, one can observe that within these ranges, the scalar field density dominates at the current epoch and stabilizes at point $P_{4}$. However, from the Hubble evolution, it can be seen that as the redshift increases ($z>0.5$), the trajectories tend to deviate from the data points. While the evolution of the distance modulus does not exhibit significant changes. To obtain the best-fit parameter values of these parameters, we will utilize Bayesian statistics as described in Sec. IV, using a similar range.
Figure 1: The evolution of dynamical variables in the quintessence model where $H_{0}=66\ \rm Kms^{-1}Mpc^{-1}$.
III K-essence field
@@
where the $k$-essence Lagrangian becomes the pressure of the field. The pressure of the $k$-essence field is assumed to have a particular form as follows:
@@
Here, $V(\phi)$ represents the potential of the field, and $F(X)$ depends on the kinetic term of the field, where $X=-\frac{1}{2}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi$. The potential extensively studied corresponding to this particular form of the Lagrangian, which exhibits scaling behavior in the past epoch and tracking behavior in the current epoch, is the inverse square type of potential given by:
$V(\phi)=\frac{\delta^{2}}{\kappa^{2}\phi^{2}}.$ (20)
The kinetic function of the field takes the form:
@@
Upon varying the action with respect to $\phi$, the field equation of motion becomes:
@@
where the subscript associated with the comma denotes the partial derivative with respect to that quantity. In the apatially flat FLRW metric, the Friedmann equation becomes:
@@
III.1 Dynamical system stability of $k$-essence field
After evaluating the background dynamical equations in the aforementioned section, we will analyze the stability of the system by defining a set of dimensionless variables as:
@@ III.1 Dynamical system stability of $k$-essence field
We choose the variables corresponding to the field sector such that the Hubble equation takes a similar form to the quintessence field. The primary variables associated with the $k$-essence field are $x,y>0$, and are assumed to be positive. Following the Hubble equation Eq. (26), the primary variables are $(x,y,\Omega_{m})$ constrained as:
@@
The second Friedmann equation can be expressed in terms of dynamical variables as:
@@
$P_{3}$ $\left(-\frac{\sqrt{3\lambda^{2}+16}+\sqrt{3}\lambda}{2\sqrt{2}},\frac{1}{2}\sqrt{3\lambda^{2}+\sqrt{9\lambda^{2}+48}\lambda+12}\right)$ $1$ 0 $\frac{1}{6}\left(-3\lambda^{2}-\sqrt{9\lambda^{2}+48}\lambda-6\right)$ Fig. [2a]
$P_{4}$ $\left(\frac{\sqrt{3\lambda^{2}+16}-\sqrt{3}\lambda}{2\sqrt{2}},\frac{1}{2}\sqrt{3\lambda^{2}-\sqrt{9\lambda^{2}+48}\lambda+12}\right)$ 1 0 $\frac{1}{6}\left(-3\lambda^{2}+\sqrt{9\lambda^{2}+48}\lambda-6\right)$ Fig. [2b]
@@
The system produces four critical points, and we have considered only those points for which $y>0$.
@@
  • Point $P_{3}$: The coordinates are $\lambda$ dependent, and the field density parameter dominates over the fluid density. The effective EoS, $\omega_{\rm eff}$ is $\lambda$ dependent and produces an accelerating solution, i.e., $-1\leq\omega_{\rm eff}\leq-1/3$, for $-0.82<\lambda<0$ and a phantom solution, $-1.2\leq\omega_{\rm eff}<-1$, for $0.1620<\lambda<$ 0.162. In this range, the real part of the eigenvalues are negative, as shown in Fig. [2a], hence the point becomes stable and shows the late-time behavior of the universe.
    Figure 2: The evolution of eigenvalues at points $P_{3,4}$ against $\lambda$ shown in Figs. [2a, 2b] respectively.
  • Point $P_{4}$: Similar to point $P_{3}$, this point produces acceleration for $0.820<\lambda<$ 0.82 and phantom-type characteristics for $-0.162<\lambda<0$. In this range, the eigenvalues yield negative non-zero real parts, as shown in Fig. [2b], indicating the stable behavior of the point. As this point also features a stable accelerating solution, it is physically viable.
@@
Based on the above critical point analysis, we see that the points $P_{3,4}$ produce accelerating and phantom solutions based on the range of $\lambda$. However, these two points do not signify the same state of the universe for a particular value of $\lambda$, neither do they exist in the same quadrant of the phase space as the coordinates become different. For example, for $\lambda=0.2$, the coordinates of $P_{3}$ become $P_{3}=(-1.54,1.84)$, while $P_{4}=(1.30,1.64)$. One of the major differences can be seen in the sign of the $x$-coordinate. So, while doing the numerical evolution of the differential equations, as stated earlier, the autonomous equations present a submanifold; consequently, the initial conditions in $x$, if chosen positive, then the system evolves and stabilizes at $P_{4}$. In terms of phase space, one can say that all the trajectories that have originated in the positive $x$-axis will remain bound in the first quadrant provided $y>0$.

### v2

# Comprehensive Study of $k$-essence Model: Dynamical System Analysis and Observational Constraints from Latest Type Ia Supernova and BAO Observations

@@ Abstract
We constrain the parameters of the $k$-essence scalar field model with inverse square and exponential potentials using data sets including Pantheon+SHOES and the Dark Energy Survey (DES) of Type Ia supernovae, Baryon Acoustic Oscillation (BAO) data from SDSS and DESI surveys, and direct measurements of the Hubble parameter and redshift obtained from the differential age method (CC). We also provide a brief perspective on the dynamical evolution of both models and derive stability constraints on the model parameters, which are then used to set appropriate priors. We adopt a Bayesian inference procedure to estimate the model parameters that best fit the data. A comprehensive analysis in light of observational data shows that the $k$-essence model fits well across all data combinations. However, according to the BIC criterion, the $\Lambda$CDM model provides a slightly better fit compared to the $k$-essence model.
@@ I Introduction
Over the past two decades, researchers have explored various alternative scenarios that can mimic the dynamics of $\Lambda$CDM. One of the most popular and heavily investigated mechanisms to tackle this problem is via the introduction of a scalar field which acts as the dark energy constituent. Inspired by particle physics, a quintessence-type field with a potential term can generate sufficient acceleration in the presence of dark matter [19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. Although the quintessence field minimally coupled with dark matter fluid can mimic the dynamical aspect of $\Lambda$CDM, it also suffers from fine-tuning of the potential parameter. Later, some string-inspired models, namely $k$-essence fields with non-linear kinetic terms, have been widely studied. It is claimed that these theories can address some of the shortcomings of the above models. Due to the associated non-linear kinetic terms, these fields have been termed non-canonical fields [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. Despite its success in explaining late-time cosmic acceleration, only a small number of attempts have been made to test the k-essence model against observational data to see how well they fit with late-time cosmological observations.
Testing the $k$-essence model against observational data is crucial to determine how well it aligns with the observations. To perform this analysis, we numerically solve the field and Friedmann equations, thereby obtaining the Hubble parameter as a function of redshift. Setting initial conditions and priors for the model parameters involves conducting a dynamical system analysis. Stability techniques within dynamical systems are invaluable tools for examining complex models, as they facilitate the identification of physically viable solutions by analyzing the existence and nature of critical points, thus constraining the model parameters [41, 42, 43, 44, 45, 46, 47, 48]. These critical points correspond to the various phases of the universe. Some of the critical points are practically useful as they may have a connection to the late-time acceleration of the universe.
To assess the stability of the critical points, we establish a system of first-order autonomous differential equations. By linearizing these autonomous equations, we can construct a Jacobian matrix and calculate the corresponding eigenvalues. If any real part of the eigenvalues is zero, the linearization method may not be applicable for determining stability. In such cases, more advanced mathematical approaches, such as the center manifold theorem, become necessary to accurately evaluate the stability.
In this study, we analyze the $k$-essence scalar field with the Lagrangian as $\mathcal{L}_{\phi}=V(\phi)(-X+X^{2})$, where $X=-\frac{1}{2}g^{\mu\nu}\nabla_{\mu}\phi\nabla_{\nu}\phi$, with two types of potentials: (i) inverse square potential $V(\phi)\propto\frac{1}{\phi^{2}}$ and (ii) exponential potential $V(\phi)\propto e^{\phi}$. We test these models using observational datasets including Cosmic Chronometer (CC), Baryonic Acoustic Oscillation (BAO), Dark Energy Survey Supernova 5YR (DES-SN), and SH0ES-calibrated Pantheon samples (Pantheon+SH0ES) of Type Ia supernovae. While extensive research has been conducted on the $k$-essence field with the inverse square potential, there is limited literature addressing the exponential potential within the Lagrangian framework specified above. Moreover, studies on pure kinetic $k$-essence fields are sparse [49, 50]. Our analysis indicates that the exponential potential model can exhibit a late-time cosmic accelerating phase, preceded by an attracting matter phase that persists for a brief period.
The data analysis reveals that both models are consistent with the observational datasets, presenting a compelling alternative to the $\Lambda$CDM model. The Akaike Information Criterion (AIC) and Bayesian Information Criterion (BIC) both strongly support the current models over $\Lambda$CDM when considering the combined datasets.
The paper is organized as follows: In Sec. II, we analyze the $k$-essence model with inverse square and exponential type potentials from a dynamical system perspective and present the numerical evolution of the cosmological parameters. The data analysis of the $k$-essence models is discussed in Sec. III. Finally, we present our conclusions in Sec. IV.
II K-essence field
@@
where the $k$-essence Lagrangian $\mathcal{L}_{\phi}$ is related to the pressure of the field, and $\mathcal{L}_{m}$ denotes the Lagrangian for the perfect fluid. In our system of units $\kappa^{2}=8\pi G$ where $G$ is Newton’s universal gravitational constant. The pressure of the $k$-essence field is assumed to have a particular form as follows:
@@
Here, $V(\phi)$ represents the potential of the field, and $F(X)$ depends on the kinetic term of the field, where $X=-\frac{1}{2}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi$. The potential extensively studied corresponding to this particular form of the Lagrangian, which exhibits scaling behavior in the past epoch and tracking behavior in the current epoch, is the inverse square type of potential [9, 33, 51]. However, in this paper our study is not limited to this particular type of potential. We will expand our study to include two types of potentials, namely the inverse square and exponential type potentials, referred to in the text as Model I and Model II. The forms of the potentials are given by:
[…]
Here, the parameters $\delta$ and $\beta$ are dimensionless constants, whereas $H_{0}$ refers to a constant whose dimension is the same as the Hubble parameter. The other parameter $V_{0}$ is a dimensional constant with mass dimension four. We have considered the exponential potential because this type of potential is usually studied with quintessential types of scalar fields [23]. However, very few studies have been conducted with the $k$-essence field [52, 53, 54], particularly with the Lagrangian form $\mathcal{L}_{\phi}=-V(\phi)F(X)$. Hence, we will carry out a detailed study of the dynamics of the $k$-essence field with both potentials. To study the following scenario, we will take the particular form of the kinetic function as given by:
@@
Here, $\rho_{m}$ and $P_{m}$ denote the background matter fluid energy density and pressure, respectively. In this paper, we choose the background fluid to be pressureless, i.e., $P_{m}=\omega_{m}\rho_{m}=0$. Upon varying the action with respect to $\phi$, the field equation of motion becomes:
@@
where the subscript associated with the comma denotes the partial derivative with respect to that quantity. In the spatially flat Friedmann-Lemaître-Robertson-Walker (FLRW) metric, $ds^{2}=-dt^{2}+a(t)^{2}d\vec{x}^{2}$, where $a$ denotes the scale factor, the Friedmann equations become:
@@
II.1 Stability of the system in the presence of Inverse square potential
After evaluating the background dynamical equations in the aforementioned section, we will analyze the stability of the system corresponding to the inverse square type of potential by defining a set of dimensionless variables as:
@@ II.1 Stability of the system in the presence of Inverse square potential
Although a large section of the literature has been published discussing the dynamics corresponding to the inverse square potential, we are explicitly showing it here for the sake of consistency. To analyze the dynamics, we choose a different dynamical setup to comprehend the dynamics of the phase space. While the dynamical variables are defined differently, the results we obtained match those of previous studies. We choose the variables corresponding to the field sector such that the Hubble equation takes a similar form to the quintessence field as presented in Ref.  [9]. The primary variables associated with the $k$-essence field are $(x,y>0)$, and are assumed to be positive. The variables $(\Omega_{\phi},\Omega_{m})$ are secondary variables as they can be expressed in terms of primary variables. The other variable, $\lambda$, is called the slope of the potential, which in this case depends on $\delta$ and becomes a constant. It should be noted that the dynamical variables defined here will not be valid for all types of potential. For instance, the dynamical variables chosen for the exponential potential in the next section in Eq. (23), although using similar notations, have definitions that are quite different from those used here. Following the Hubble equation Eq. (9), the primary variables are $(x,y,\Omega_{m})$ constrained as:
@@
Similarly, the second Friedmann equation can be expressed in terms of dynamical variables as:
@@
$P_{3}$ $\left(-\frac{\sqrt{3\lambda^{2}+16}+\sqrt{3}\lambda}{2\sqrt{2}},\frac{1}{2}\sqrt{3\lambda^{2}+\sqrt{9\lambda^{2}+48}\lambda+12}\right)$ $1$ 0 $\frac{1}{6}\left(-3\lambda^{2}-\sqrt{9\lambda^{2}+48}\lambda-6\right)$ Fig. [1a]
$P_{4}$ $\left(\frac{\sqrt{3\lambda^{2}+16}-\sqrt{3}\lambda}{2\sqrt{2}},\frac{1}{2}\sqrt{3\lambda^{2}-\sqrt{9\lambda^{2}+48}\lambda+12}\right)$ 1 0 $\frac{1}{6}\left(-3\lambda^{2}+\sqrt{9\lambda^{2}+48}\lambda-6\right)$ Fig. [1b]
@@
The system produces four critical points, and we have considered only those points for which $y>0$. As the matter sector is assumed to be made up of dark matter, our assumption implies that the fluid representing dark matter is pressure-less $\omega_{m}=0$. We use linearization techniques to determine the stability of the critical points by evaluating the Jacobian matrix defined as follows:
$J_{ij}=\frac{\partial x_{i}^{\prime}}{\partial x_{j}}\bigg{|}_{\bm{x}_{*}},$ (17)
where $x_{i}=\{x,y,...\}$ for $i=1,2,...$ and $\bm{x}_{*}$ represent the set of coordinates specifying a critical point. Corresponding to this autonomous system, the Jacobian matrix becomes $n\times n$ dimensional. To assess stability, we determine the eigenvalues and based on the signature of the real parts of eigenvalues, the point can be categorized as stable, unstable, or saddle. If all the real parts of the eigenvalues are negative (positive), the corresponding point becomes asymptotically stable (unstable). However, for alternating signs, the point is labeled as a saddle. The detailed description of the critical points and corresponding stability is discussed as follows:
@@
  • Point $P_{3}$: The coordinates are $\lambda$ dependent, and the field density parameter dominates over the fluid density. The effective EoS, $\omega_{\rm eff}$ is $\lambda$ dependent and produces an accelerating solution, i.e., $-1\leq\omega_{\rm eff}\leq-1/3$, for $-0.82<\lambda<0$ and a phantom solution, $-1.2\leq\omega_{\rm eff}<-1$, for $0.1620<\lambda<$ 0.162. In this range, the real part of the eigenvalues are negative, as shown in Fig. [1a], hence the point becomes stable and shows the late-time behavior of the universe.
    Figure 1: The evolution of eigenvalues at points $P_{3,4}$ against $\lambda$ shown in Figs. [1a, 1b] respectively.
  • Point $P_{4}$: Similar to point $P_{3}$, this point produces acceleration for $0.820<\lambda<$ 0.82 and phantom-type characteristics for $-0.162<\lambda<0$. In this range, the eigenvalues yield negative non-zero real parts, as shown in Fig. [1b], indicating the stable behavior of the point. As this point also features a stable accelerating solution, it is physically viable.
@@
Based on the above critical point analysis, we see that the points $P_{3,4}$ produce accelerating and phantom solutions based on the range of $\lambda$. However, these two points do not signify the same state of the universe for a particular value of $\lambda$, neither do they exist in the same quadrant of the phase space as the coordinates become different. For example, for $\lambda=0.2$, the coordinates of $P_{3}$ become $P_{3}=(-1.54,1.84)$, while $P_{4}=(1.30,1.64)$. One of the major differences can be seen in the sign of the $x$-coordinate. Due to the existence of an invariant submanifold, the initial conditions in $x$, if chosen positive, the system evolves and stabilizes at $P_{4}$. In terms of phase space, one can say that all the trajectories that have originated in the positive $x$-axis will remain bound in the first quadrant provided $y>0$.